Shift Summary for EVENING shift, 25 July 99

Arun Tripathi (arun@physics.ucla.edu)
Sun, 25 Jul 1999 22:02:02 -0700 (PDT)

Shift Summary for EVENING shift, 25 July 99

On Shift: A. Soddu, A. Tripathi

Hours of beam delivered at all intensities: ~7.5
Hours of beam delivered at nominal intensity: ~7 (@3-3.5E12)
Hours of E832 data (special runs not included): 5 hr 46 min
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
KTEV2 crash, stopping a hung run, changing tapes, drt_fresh_starts.

========================================================================
Runs ended this shift:

14021 --- BEAM_832 (BOA >2.4)

Tapes brought to FCC this shift: None

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
Inherited a bad-beam (BOA >2.4) run, KTEV2 died and was rebooted.
Good (although low int., SEM ~3E12, inches up to 3.5E12 around spill 115,
run 14023) beam since KTEV2 was resurrected. Pretty smooth and quiet
running since then.

========================================================================
Description of Shift:

15:30 Inherited a run with bad beam quality. BOA ~2.4. Day people told
MCR to fix it, and they are working on it.

16:09 Paged Rick Coleman about the BOA. He promises to talk to the
beam people about it.

16:39 Rick C. called back. He spoke with MCR, and they are still tuning
the beam, will call experts.

17:09 DDD Veto 2 enabled because KTEV2 crashed (took us a while to figure
out that's the reason).

We find out that the pipe.tcl program doesn't work, it just pops
up a blank window.

Tried stopping run 14021, but it hanged. After saving the online
histos, killed the run (drt_fresh_start -y). The tapes were
70% full, and upon Rick's suggestion, they were taken out and
fresh tapes loaded into the drives.

MCR calls and tells us beam quality has improved. BOA is now
~1.25.
18:01 KTEV2 rebooted with Ed's help (on the phone).

18:23 Beam 832 run 14023 started.

19:47 SEM has been low for this run, around 3E12. Called MCR. They are
aware of it and working on it.

19:54 Restarted processes "check_scalers" and "mon_memory" on ktev3
in order to convert the red "SADQ" button back to green.

20:54 Getting aborts. No beam for 10 minutes.
Even during the begining of the run, there were several aborts
and SY conservation periods.

~21:00 SEM is up to 3.5E12. MCR is still "tuning for higher intensity".

23:58 Loaded onsplit tape QKNL14 in drive tps3d2 on ktev4.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

No muon run or change of magnet polarity tonight. Continue to take
BEAM_832 data under the same conditions.